# Short-Range Order Based Ultra Fast Large-Scale Modeling of High-Entropy Alloys

### v2

@@ 3 Discussion
This study demonstrates PyHEA’s transformative impact on high-entropy alloy modeling through three key achievements. First, PyHEA achieves unprecedented computational efficiency while maintaining high accuracy. Our benchmarks show that PyHEA outperforms traditional tools like ATAT and SCRAPs by several orders of magnitude, delivering speedups of up to 333,000× and 13,900× respectively (Table 1). This dramatic improvement in performance does not come at the cost of accuracy—PyHEA consistently produces more stable and accurate results across various system sizes, as evidenced by the tighter fitness distributions shown in Fig. 2.
Second, PyHEA successfully addresses the scalability challenge that has long constrained HEA modeling. While existing methods become computationally prohibitive beyond a few thousand atoms, PyHEA’s GPU-accelerated framework efficiently handles systems exceeding 256,000 atoms. This capability represents a significant breakthrough, enabling simulations at experimentally relevant scales that were previously inaccessible. The practical impact of this scalability is demonstrated in our Fe-Mn-Cr-Co system simulation, where PyHEA reduced computation time from days to mere seconds while accurately reproducing temperature-dependent SRO patterns documented in previous studies.
Third, PyHEA’s innovative approach to SRO calculations through incremental updates and GPU optimization establishes a new standard for computational efficiency in materials modeling. The incremental SRO algorithm reduces computational overhead by up to 99.6% (Fig. 5), while GPU acceleration delivers over 1,000× speedup compared to CPU-based calculations. These technical advances make it feasible to conduct large-scale atomistic simulations with unprecedented efficiency, opening new possibilities for materials design and optimization.
The broader implications of PyHEA’s capabilities extend beyond pure computational advancement. The dramatic reduction in simulation time from days to minutes facilitates more iterative and exploratory research approaches. Researchers can now rapidly test hypotheses, adjust parameters, and explore diverse atomic configurations that were previously impractical due to computational constraints. This acceleration of the research cycle has the potential to significantly accelerate the discovery and optimization of new high-entropy alloys.
Furthermore, PyHEA’s success in modeling complex SRO phenomena in large-scale systems provides valuable insights for understanding structure-property relationships in HEAs. The ability to accurately simulate and analyze SRO patterns across different temperatures and compositions, as demonstrated in our Fe-Mn-Cr-Co case study, offers a powerful tool for predicting and optimizing material properties. This capability is particularly crucial given the growing recognition of SRO’s role in determining mechanical and thermal properties of HEAs. Future developments could further enhance PyHEA’s capabilities in several directions. Integration with advanced machine learning models, particularly reinforcement learning approaches, could optimize the search for desired atomic configurations. Expanding the framework to incorporate additional physical properties and interaction models would broaden its applicability across materials science. Additionally, collaborative efforts with experimental teams could strengthen the validation of PyHEA’s predictions under various thermomechanical conditions.
In conclusion, PyHEA represents a significant advancement in computational materials science, bridging the gap between theoretical understanding and practical simulation of large-scale HEA systems. Its combination of unprecedented computational efficiency, accurate SRO modeling, and scalability to experimentally relevant sizes establishes a new paradigm for materials design and optimization. As the field continues to evolve, PyHEA’s framework provides a robust foundation for future innovations in high-entropy alloy research and development.
@@ 4 Methods
Figure 4: Workflow of the PyHEA algorithm for optimizing atomic configurations. (a) The global search framework iteratively cycles through the stages of perturbation, local search, and ranking to converge toward an optimal solution. (b) Initial atomic configurations consist of randomly generated and previously identified optimal structures for exploration. (c) Monte Carlo (MC) methods are applied to refine each candidate configuration and evaluate their fitness. (d) Configurations are ranked based on fitness, and the lowest-ranked fraction of solutions is discarded to enhance the selection process for subsequent iterations.
@@ 4.1 Global Search
1:Fix input and optimization function
2:Optimized solution
3:Initialize solutions
@@
5:     Create new solutions randomly
6:     Calculate fitness $F$ of the solutions
7:     Choose fraction of solutions with best fitness (top solutions)
8:     for each solution $\in$ top solutions do
9:         Search using Monte Carlo (Local Search)
@@
11:     Discard fraction $p_{a}$ of worst solutions
12:     Rank the solutions and find the current best
@@
14:Return the best solutions
The optimization of atomic configurations in High-Entropy Alloys (HEAs) presents a complex challenge due to the vast solution space. As illustrated by Fig. 4, our global search framework addresses this challenge through an iterative approach that combines systematic exploration with targeted refinement.
As shown in Fig. 4(a), the global search phase is essential for navigating the extensive configuration space of atomic arrangements in HEAs. This phase strikes a balance between exploration and refinement, allowing for the identification of promising regions within the solution space.
The search process begins by generating candidate solutions that represent different atomic configurations (Fig. 4(b)). During each iteration, a subset of solutions undergoes random perturbations to enable broader exploration, while the top-performing configurations are preserved for further refinement (Fig. 4(d)). This approach effectively integrates random exploration with solution improvement.
Drawing inspiration from the Cuckoo Search algorithm, the method employs L’evy flights for optimization [36, 23]. These random jumps facilitate global exploration and help prevent the algorithm from becoming trapped in local minima. Each candidate solution is evaluated using a fitness function, $F(X)$, which quantifies the quality of SRO in the atomic structure by comparing local atomic arrangements to the desired SRO characteristics.
To maintain solution diversity, a portion of low-performing solutions is eliminated at the end of each iteration, while the best solutions are carried forward to the next cycle. This iterative procedure continues until convergence, gradually driving the system toward an optimized atomic configuration with enhanced SRO properties.
@@ 4.2 Local Search
1:Fix input and optimization function
2:Optimized solution
3:Initialize current configuration
4:while iteration $<$ Local iterations do
5:     Perturb configuration randomly
6:     Calculate fitness difference $\delta F=F(x)-F(x_{\text{new}})$
@@
8:         Restore fitness and configuration to previous state
@@
11:Return the best solutions
@@
1:Fix input and optimization function
2:Optimized solution
3:Initialize current configuration
@@
5:     Load current fitness and configuration
6:     Perturb configuration randomly
7:     Calculate fitness difference $\delta F_{i}=F(x)-F(x_{i})$
8:     Synchronize tasks and sort $\delta F_{i}$ in descending order
9:     if $\delta F_{min}>0$ or $|\delta F_{min}|\leq\text{threshold}$ then
10:         $depth$ = 0
11:         Update current fitness and configuration to $x_{i}$
@@
19:Return the best solutions
The local search phase plays a critical role in refining candidate solutions obtained from the global search phase, particularly in large-scale HEA simulations. To achieve high performance and precision, a parallelized Monte Carlo (MC) search (Alg. 3)is employed, leveraging GPU implementation to efficiently refine atomic configurations in parallel. This parallelization enables the distribution of tasks across multiple threads, where configurations are perturbed, fitness differences are evaluated, and only the highest-performing solutions are retained. This method enhances both computational speed and accuracy, significantly reducing the overall computation time.
Compared to the serial version (Alg. 2), which processes configurations one at a time and struggles with the scalability required for large systems, the parallel approach addresses this limitation by processing multiple configurations simultaneously. This parallelization significantly enhances both scalability and performance. By integrating the parallelized local search with the global search phase, candidate solutions are refined to achieve optimal local configurations, facilitating convergence to high-quality results with improved SRO properties.
@@ 4.3 Short-Range Order
Short-Range Order (SRO) characterizes the local atomic structure in HEAs, describing the tendency of certain atomic species to preferentially cluster compared to a random distribution [37]. The SRO is typically quantified by the Warren-Cowley parameter [9], denoted as $\alpha_{ij}^{k}$, which measures the degree of ordering between atomic species $i$ and $j$. For a given atom $i$, the SRO parameter for its $k$-th coordination shell is defined as:
$\alpha_{ij}^{k}=1-\frac{P_{ij}^{k}}{c_{j}},$ (2)
where $P_{ij}^{(k)}$ is the conditional probability of finding a $j$ atom in the $k$-th coordination shell of an $i$ atom, $c_{j}$ is the overall concentration of $j$ atoms in the alloy.
The SRO parameter $\alpha_{ij}^{k}$ ranges from $[-1,1]$:
@@
    $\alpha_{ij}^{k}<0$: Indicates an attraction between elements $i$ and $j$, suggesting a preference to be adjacent.
@@
    $\alpha_{ij}^{k}>0$: Indicates a repulsion between elements $i$ and $j$, suggesting a preference to be separated.
For each atomic pair type $(i,j)$, the SRO parameter for the entire system of $N$ atoms can be expressed as:
$\langle\alpha_{ij}^{k}\rangle=\frac{1}{N_{i}}\sum_{n=1}^{N_{i}}\left(1-\frac{P_{ij}^{k}}{c_{j}}\right),$ (3)
where $N_{i}$ is the number of atoms of type $i$.
To calculate this parameter, the number of $j$ atoms within each coordination shell around every $i$ atom is determined, and the result is normalized by the total concentration of $j$ atoms ($c_{j}$). computed SRO values for each pair $(i,j)$ are then averaged over the entire system to obtain a global measure of the SRO.
4.4 Molecular Dynamics
Molecular dynamics (MD) simulations were performed using LAMMPS [38, 39, 40], employing the 2-NN MEAM potential for the Fe-Mn-Cr-Co alloy system. The simulations encompassed system sizes ranging from 256 to 256,000 atoms, with a timestep of 2 fs in an NPT ensemble. The built-in MC method in LAMMPS was utilized to compute the Radial Distribution Function (RDF), providing detailed insight into the atomic arrangement and SRO.
These simulation parameters were chosen to ensure reproducibility and to facilitate replication in future studies. The results of these MD simulations are discussed in detail in section 2, demonstrating the effectiveness of our method.
4.5 Incremental Calculation of SRO (Fitness)
In the Monte Carlo (MC) search process, the SRO parameter is crucial for evaluating the fitness of each atomic configuration. As previously discussed, SRO quantifies the local atomic arrangement, guiding the search toward optimal configurations during each MC iteration.
In the traditional non-incremental approach, SRO is recalculated from scratch at every MC iteration (Equation 3). This recalculation becomes increasingly computationally expensive as system size grows. Fig. 5 illustrates that for larger systems, the SRO calculation can consume more than 99% of the total computational cost, highlighting the necessity for a more efficient method.
To address this challenge, we implemented an incremental SRO calculation. Instead of recalculating the SRO for the entire system at each iteration, we update only the SRO values influenced by the random swap of two atoms, $i$ and $j$. This approach localizes the update to the environments of the swapped atoms and their neighbors, enabling a more efficient adjustment of the SRO during each MC iteration.
For the incremental update, we first compute the change in SRO for the atoms directly involved in the swap. The change in the SRO value for atom $i$ and its neighbors is given by:
$\Delta\alpha_{ij}^{k}=\left(1-\frac{P_{ij}^{k}\text{(new)}}{c_{j}}\right)-\left(1-\frac{P_{ij}^{k}\text{(old)}}{c_{j}}\right),$ (4)
[…]
Benchmark data shown in Fig. 5 highlights the significant computational burden of non-incremental SRO calculations, which account for 97.4% to 99.6% of the total execution time as system sizes increase from 432 to 3,456 atoms. However, the incremental SRO approach substantially alleviates this burden, reducing total execution times to 71.8% for 432 atoms and stabilizing around 65.5% for larger systems. These findings underscore the incremental method’s ability to enhance computational efficiency and scalability for large-scale HEA simulations.
By adopting incremental SRO updates, we maintain the accuracy of fitness evaluations while significantly reducing the computation time for each MC iteration. This ensures that the MC search process remains efficient and scalable for large-scale HEA simulations.
4.6 GPU Optimization
To further accelerate the MC search and enhance scalability, we implemented a parallel GPU-based optimization. Each thread processes an independent task, and each thread block manages the complete MC search for a single solution. This approach allows the GPU to handle multiple solutions simultaneously, leveraging resources such as registers, shared memory, and low-precision data formats to maximize computational efficiency.
4.6.1 Parallel Processing on the GPU
In this setup, each thread performs the following tasks:

### v3

@@ 3 Discussion
This study presents PyHEA as a computational framework that significantly advances HEA modeling by offering dramatic improvements in simulation speed without compromising accuracy. Compared to established tools such as ATAT-mcsqs and SCRAPs, PyHEA achieves speedups of up to 333,000× and 13,900×, respectively (Table 1), while consistently demonstrating tighter fitness distributions (see Fig. 2). These findings confirm that PyHEA maintains or even enhances precision across a wide range of system sizes.
A key factor in PyHEA’s scalability is its ability to handle systems with more than 256,000 atoms. Previous approaches often become intractable beyond a few thousand atoms, yet PyHEA’s GPU-accelerated design and incremental fitness calculations (detailed in subsection 4.5) make large scale simulations both feasible and efficient. In the Fe–Mn–Cr–Co case study, for instance, PyHEA reduces computation time from several hours or days to minutes while reliably reproducing the temperature-dependent SRO patterns demonstrated in previous work.
To explain why PyHEA outperforms existing methods such as SCRAPs, it is crucial to highlight two technical innovations. First, the incremental SRO algorithm recalculates only the local environment affected by each atomic swap rather than recomputing the entire system’s SRO at every iteration. As described in Section 4.5, this approach lowers SRO-related overhead by up to 99.6%. Second, PyHEA capitalizes on GPU acceleration to parallelize local Monte Carlo perturbations and incremental fitness evaluations, thereby multiplying the benefits of the incremental SRO strategy. In contrast, SCRAPs relies on a more conventional, system-wide recalculation of SRO and does not exploit large-scale GPU parallelization; hence it cannot match PyHEA’s order-of-magnitude speedups.
Beyond delivering efficiency and accuracy, PyHEA significantly accelerates the research cycle itself. By reducing simulation time from days to mere minutes or seconds, researchers can now quickly test hypotheses, tune parameters, and explore diverse atomic configurations that were previously too time-consuming. This capability supports a more iterative approach to materials design, enabling deeper and faster insights into composition–structure–property relationships in HEAs.
Moreover, PyHEA’s ability to simulate large systems with complex SRO phenomena under various temperatures and compositions provides valuable information for understanding structure-sensitive properties. In particular, it opens opportunities for investigating how SRO influences mechanical and thermal behaviors. The demonstrated consistency with both experimental and prior computational data reinforces the reliability of PyHEA’s predictions in practical contexts.
Looking ahead, PyHEA can be extended through integration with machine learning models, especially reinforcement learning, to guide or refine its global search. Incorporating more sophisticated interatomic potentials or linking to first-principles simulations would further broaden its applicability. Additional validation in collaboration with experimental teams would strengthen confidence in PyHEA’s predictive accuracy under different thermomechanical conditions.
In conclusion, PyHEA represents a substantial breakthrough in computational materials science by bridging the gap between theory and large-scale HEA simulations. Its use of incremental SRO calculations and GPU acceleration explains its marked speedups over methods like SCRAPs. By enabling full-atomistic simulations of tens or hundreds of thousands of atoms in a fraction of the time, PyHEA paves the way for faster, more exploratory HEA research and optimization, thus laying a foundation for future advancements in alloy development and discovery.
@@ 4 Methods
This section details the methods employed in our framework for optimizing atomic configurations in HEAs. We discuss the global search at subsection 4.1 for exploring the configuration space, the local search at subsection 4.2 for refining promising candidates, SRO metrics at subsection 4.3 for quantifying atomic arrangements, fitness calculation at subsection 4.4 for evaluating configurations, incremental updates at subsection 4.5 for enhancing computational efficiency, GPU implementation at subsection 4.6 for boosting the performance with heterogeneous computing devices, and finally MD simulations at subsection 4.7 for structure optimization and analysis. Fig. 5 provides an overview of the PyHEA workflow, summarizing its main components and their interactions.
Figure 5: Workflow of the PyHEA algorithm for optimizing atomic configurations. (a) The global search framework iteratively cycles through the stages of perturbation, local search, and ranking to converge toward an optimal solution. (b) Initial atomic configurations consist of randomly generated and previously identified optimal structures for exploration. (c) Monte Carlo (MC) methods are applied to refine each candidate configuration and evaluate their fitness. (d) Configurations are ranked based on fitness, and the lowest-ranked fraction of solutions is discarded to enhance the selection process in subsequent iterations.
@@ 4.1 Global Search
1:Input parameters, optimization function
2:Optimized solution(s)
3:Initialize solutions (candidate configurations)
@@
5:     Create new solutions via random perturbations
6:     Calculate fitness $F$ for all solutions
7:     Select a fraction of solutions with the best fitness (top solutions)
8:     for each solution in top solutions do
9:         Refine solution via Monte Carlo (local search)
@@
11:     Discard the worst fraction $p_{a}$ of solutions
12:     Rank the solutions and identify the current best
@@
14:return best solutions
The optimization of atomic configurations in HEAs presents a complex challenge due to the vast solution space. As illustrated in Fig. 5(a), our global search framework implements an iterative optimization strategy that progressively refines atomic configurations through multiple cycles of exploration and evaluation, following the main workflow outlined in Algorithm 1. The effectiveness of this strategy is governed by several key parameters, including the number of candidate solutions and global iterations, which significantly influence the balance between exploration breadth and computational efficiency (as detailed in  subsection 2.1).
The search process begins with the initialization of candidate solutions representing different atomic arrangements (Fig. 5(b)). The number of these initial solutions directly affects the search space coverage, with larger solution pools enabling more comprehensive exploration at the cost of increased computational overhead. During the first iteration, these configurations are generated entirely randomly to ensure comprehensive coverage of the solution space. In subsequent iterations, the solution pool combines newly generated random configurations with top-performing solutions retained from the previous iteration. Crucially, we maintain the number of retained solutions at less than or equal to half of the total configurations selected for local search, thereby ensuring sufficient randomness in each iteration while preventing premature convergence to local optima.
Each candidate solution undergoes fitness evaluation (subsection 4.4) to quantify its alignment with target SRO characteristics. Following evaluation, a predetermined number of the highest-ranking configurations are selected for refinement through local search (Fig. 5(c)). The number of parallel tasks during this phase and the search depth of the Monte Carlo optimization significantly impact both the quality of refinement and computational efficiency. This local optimization phase, detailed in  subsection 4.2, employs Monte Carlo perturbations to systematically explore the neighborhood of promising configurations, thereby improving their SRO properties.
The refined configurations then undergo a comprehensive ranking process, as illustrated by Fig. 5(d). At each iteration’s conclusion, a fixed fraction $p_{a}$ of the lowest-performing solutions is eliminated, while the top-performing configurations are preserved as seeds for the subsequent iteration. This selection mechanism, combined with the continuous introduction of new random configurations, maintains an effective balance between preserving successful structural patterns and exploring new possibilities within the vast configuration space.
This iterative optimization process, as described in Algorithm 1, continues until reaching the maximum number of global iterations. The optimal number of iterations, as demonstrated in  subsection 2.1, is crucial for achieving convergence while maintaining computational efficiency. Through careful parameter tuning, the algorithm progressively converges toward optimal atomic configurations that exhibit desired SRO characteristics.
@@ 4.2 Local Search
1:Configuration $x$, optimization function $F$
2:Refined configuration
3:Initialize current configuration $x_{\text{curr}}\leftarrow x$
4:while iteration $<$ Local_iterations do
5:     Generate a new solution $x_{\text{new}}$
6:     Calculate $\delta F=F(x_{\text{curr}})-F(x_{\text{new}})$
@@
8:         Update $x_{\text{curr}}\leftarrow x_{\text{new}}$
@@
11:return $x_{\text{curr}}$
@@
1:Configuration $x$, optimization function $F$
2:Refined configuration
3:Initialize current configuration $x_{\text{curr}}\leftarrow x$
@@
5:     Load $x_{\text{curr}}$ and $F(x_{\text{curr}})$
6:     Generate a new solution $x_{i}$
7:     Calculate $\delta F_{i}=F(x_{\text{curr}})-F(x_{i})$
8:     Synchronize tasks, then sort all $\delta F_{i}$ values
9:     if $\delta F_{\min}>0$ or $|\delta F_{\min}|\leq\text{threshold}$ then
10:         $depth\leftarrow 0$
11:         Update $x_{\text{curr}}\leftarrow x_{i}$
@@
19:return $x_{\text{curr}}$
While the global search framework effectively identifies promising configurations through iterative exploration, achieving optimal SRO characteristics requires more focused refinement. To address this need, we implement a specialized local search phase employing Monte Carlo techniques with both serial (Algorithm 2) and parallel (Algorithm 3) approaches, with the latter specifically optimized for GPU acceleration to handle large-scale systems efficiently.
In the serial implementation, each configuration undergoes sequential refinement through iterative solution generation and evaluation. The algorithm computes the change in fitness ($\delta F$) for each new solution, accepting modifications that either improve the configuration ($\delta F<0$) or meet specific threshold criteria. While this approach is effective for smaller systems, it becomes computationally intensive as system size increases, motivating our development of a parallel implementation.
The parallel Monte Carlo search (Algorithm 3) significantly enhances efficiency by simultaneously evaluating multiple candidate solutions. For each solution, multiple tasks operate in parallel, where each task generates a new configuration by randomly swapping the positions of two atoms. As demonstrated in subsection 2.1, the number of parallel tasks represents a crucial parameter, with our analysis revealing optimal performance in the range of 64–128 tasks. After each round of atomic swaps, all tasks synchronize to evaluate the fitness changes and identify the most promising configurations. This coordinated approach ensures that computational resources are focused on the most promising regions of the solution space, while the depth-based termination criterion prevents excessive computation in regions unlikely to yield significant improvements.
A key innovation in our methods is the introduction of the search depth parameter, which serves as both a convergence criterion and an efficiency control mechanism. As shown in Algorithm 3, the depth counter ($depth$) tracks consecutive non-improving iterations, incrementing when no better solutions are found and resetting to zero when improvements occur. Our parametric studies (subsection 2.1) indicate that a search depth threshold of 8–10 provides an optimal balance between solution quality and computational efficiency. When the depth counter reaches this threshold, indicating a sustained lack of improvement, the local search terminates and returns the best configuration found.
Another key innovation in our local search is the incremental fitness calculation. Instead of recalculating the entire system’s SRO after each atomic swap, we update only the local regions affected by the swap operation. For large systems containing thousands of atoms, this approach reduces the computational overhead by up to 99.6% compared to traditional full recalculation methods. The technical details and implementation of this incremental strategy are discussed thoroughly in subsection 4.5.
@@ 4.3 Short-Range Order
Short-range order (SRO) describes how atoms in a multi-component system tend to cluster or repel each other relative to a random distribution [37]. We quantify SRO using the Warren–Cowley parameter [9], denoted as $\alpha_{ij}^{(k)}$, which characterizes the degree of ordering between atomic species $i$ and $j$ in the $k$-th coordination shell:
$\alpha_{ij}^{(k)}=1-\frac{P_{ij}^{(k)}}{c_{j}},$ (2)
where $P_{ij}^{(k)}$ represents the conditional probability of finding a type-$j$ atom in the $k$-th shell around a type-$i$ atom, and $c_{j}$ is the overall concentration of type-$j$ atoms in the system. The parameter $\alpha_{ij}^{(k)}$ ranges from $-1$ to $+1$, where negative values indicate preferential attraction between species $i$ and $j$, positive values suggest repulsion, and zero corresponds to a random distribution.
4.4 Fitness Calculation
To evaluate how well a given atomic configuration matches a target SRO pattern, we define a fitness function that incorporates the Warren–Cowley parameters:
$F(X)$ $=\sqrt{\sum_{k=1}^{N_{\mathrm{shells}}}w_{k}\sum_{i=1}^{N_{\mathrm{types}}}\sum_{j=1}^{N_{\mathrm{types}}}\bigl{(}\alpha_{ij}^{(k)}(X)-\alpha_{ij}^{(k),\mathrm{target}}\bigr{)}^{2}}\,,$ (3)
where $k$ indexes the coordination shells ($1\leq k\leq N_{\mathrm{shells}}$), and $i,j$ enumerate the atomic species ($1\leq i,j\leq N_{\mathrm{types}}$). Here, $\alpha_{ij}^{(k)}(X)$ is computed according to Eq.(1) for configuration $X$. The shell-specific weights $w_{k}$ and target SRO values $\alpha_{ij}^{(k),\mathrm{target}}$ are user-specified input parameters, allowing flexible control over the optimization process. The weights $w_{k}$ can be adjusted to emphasize certain coordination shells, while $\alpha_{ij}^{(k),\mathrm{target}}$ values can be derived from experimental measurements, theoretical calculations, or other simulation results to guide the structure generation toward desired atomic ordering patterns.
4.5 Incremental Calculation of Fitness
In a non-incremental approach, the Warren–Cowley parameters and resulting fitness are recomputed from scratch at each Monte Carlo step (Equation 2 and Equation 3), which becomes computationally prohibitive for large systems. As shown in Fig. 6, for supercells containing thousands of atoms, these calculations can consume over 99% of the total execution time.
To address this bottleneck, we implement an incremental fitness update strategy. Rather than recalculating all $\alpha_{ij}^{(k)}$ values after each atomic swap, we update only those parameters affected by the exchanged atoms $i$ and $j$. The change in the Warren–Cowley parameter for the $k$-th shell is given by:
$\Delta\alpha_{ij}^{(k)}=\alpha_{ij}^{(k),\mathrm{new}}-\alpha_{ij}^{(k),\mathrm{old}}=\left(1-\frac{P_{ij}^{(k),\mathrm{new}}}{c_{j}}\right)-\left(1-\frac{P_{ij}^{(k),\mathrm{old}}}{c_{j}}\right),$ (4)
where $P_{ij}^{(k),\mathrm{new}}$ and $P_{ij}^{(k),\mathrm{old}}$ are the probabilities before and after the swap. Since only atoms within the local environment of sites $i$ and $j$ are affected, this reduces the computational complexity from $O(N)$ to $O(1)$ per Monte Carlo step.
4.6 GPU Optimization
To maximize computational efficiency for large-scale systems, we implement a parallel GPU-accelerated version of the Monte Carlo search. The implementation follows a hierarchical parallelization strategy where each GPU thread processes an individual atomic swap operation, while thread blocks manage complete Monte Carlo refinements for distinct candidate solutions. This design enables concurrent optimization of multiple configurations while efficiently utilizing GPU resources including registers, shared memory, and mixed-precision arithmetic.
4.6.1 Parallel Processing Architecture
The GPU implementation distributes computation across multiple levels:
@@ 4.6.1 Parallel Processing Architecture
    Thread-Level Parallelism: Individual threads handle SRO calculations for specific atomic pairs using reduced precision (FP16, Uint16) data types, balancing numerical accuracy with memory efficiency.
@@
    Block-Level Management: Each thread block maintains a candidate configuration in shared memory, following the lattice structure described in [41]. This arrangement minimizes global memory access during atomic position updates.
4.6.2 Incremental Fitness Evaluation
Following the incremental update strategy detailed in subsection 4.5, each thread computes local fitness changes according to:
$F_{\text{new}}=F_{\text{prev}}+\Delta F_{\text{perturb}},$ (5)
where $\Delta F_{\text{perturb}}$ is calculated from the changes in Warren–Cowley parameters ($\Delta\alpha_{ij}^{(k)}$) affected by the atomic swap. This approach maintains consistency with the fitness function defined in Equation 3 while exploiting the locality of updates.
4.6.3 Parallel Selection Strategy
After computing fitness changes, threads within each block cooperatively sort the results using shared memory operations. The best configuration update is identified according to:
$F_{\text{best}}=\min_{\text{updates}}{F_{\text{new}}},$ (6)
ensuring that each block maintains its optimal configuration throughout the refinement process.
4.6.4 Convergence Criteria
The GPU-accelerated search terminates when either: